OpenStudy (anonymous):

After 8 months, $750 simple interest was owed ona loan of $9000. Find the annual interest rate.

OpenStudy (anonymous):

750/9000 turns it into a percentage... see where that's going?

OpenStudy (anonymous):

then adjust for the portion of the year...

OpenStudy (anonymous):

or 750 / 8 times 12 all divided by 9000 is your annual rate.

OpenStudy (anonymous):

i = Prt 750 = 9000*r*(8/12) 75 = 900*r* (2/3) 75 = 600r r = 0.1250 or 12.5%

OpenStudy (anonymous):

Just like quin21 shows it using the textbook formula, solved for r.